A Fahrenheit and a centigrade thermometer are both immersed in a fluid. The Fahrenheit numerical reading is twice that of the centigrade reading. What is the temperature of the fluid in Rankine and Kelvin scale?

1
Expert's answer
2021-11-09T10:44:44-0500

Say, the temp. is T = t deg C

Therefore, F = 2t deg F


Now, "F = \frac{9}{5}C + 32"

"2t = \frac{9t}{5}+32"

10t = 9t + 160

Or t = 160 deg C = 160 + 273 = 433 deg. K

Hence T = 320 deg. F


deg Rankine = (deg. C + 273.15)*1.8 = (160 + 273.15)*1.8 = 779.67 deg R


Final Answer: Temp. T = 160 deg C, 320 deg. F, 779.67 deg. R